# Zero-Shot Video Editing through Adaptive Sliding Score Distillation

### v1

@@ 3.4 Joint Guidance Incorporating Stable Diffusion
Figure 4: Qualitative Comparison. Comparison of our ASSD with five baseline methods. Our method achieves a good balance between editing and preserving unedited information in the original video. Additionally, it possesses a certain capability for geometric shape editing.
4 Experiments
4.1 Experiments Setting
We collected approximately 30 text-video pairs from public datasets such as DAVIS [19], WebVid [2], and YouTube [1], and extracted 8 or 16 evenly spaced frames from each video for our evaluation data. We used ModelscopeT2V and Zeroscope as our base models. For optimizing the latent code of a video, we employed the SGD optimizer with a learning rate of 0.1, running the optimization for 200 steps, the process takes approximately 3 to 4 minutes on a single V100 GPU.
We evaluated our method alongside five high-performing and representative video editing methods. Tune-A-Video (TAV) [34] fine-tunes the denoising U-Net on a single video to learn the motion patterns of the video to be edited. FateZero [22] achieves zero-shot video editing through DDIM Inversion and attention manipulation techniques. FLATTEN [5] utilizes an external optical flow detection model to enable more accurate attention manipulation. TokenFlow [9] ensures the consistency of the edited video by modifying keyframes and propagating these modifications throughout the entire video. ControlVideo [37] utilizes information such as depth and edges from the original video to control the editing results.
4.2 Qualitative Experiment
Fig. 4 presents the qualitative comparison results between our method and other SOTA baselines. Our method achieves comparable or superior performance in terms of alignment with the target text prompts, preservation of the original video’s layout, motion patterns, and preservation of unedited regions.
@@ 4.2 Qualitative Experiment
TAV [34] 0.2988 0.9086 19.06 2.64 4.25
FateZero [22] 0.2738 0.9324 20.63 26.17 30.62
ControlVideo [37] 0.2916 0.9301 8.48 5.30 3.70
FLATTEN [5] 0.2931 0.9249 14.82 17.99 11.11
TokenFlow [9] 0.2910 0.9561 12.67 20.11 24.40
@@
4.3 Quantitative Experiment
@@
We utilized the CLIP model [23] for automatic evaluation, employing two metrics: CLIP Text Alignment Score and CLIP Frame Consistency Score. CLIP Text Alignment Score is defined as the mean cosine similarity between the CLIP image embeddings of the frames in the edited video and the CLIP text embedding of the target text prompt. This metric measures how well the edited video aligns with the target text prompt. CLIP Frame Consistency Score is defined as the mean cosine similarity between the CLIP image embeddings of frames within the video. This metric assesses the inter-frame consistency of the edited video. The results of the automatic metric are reported in Tab. 1.
@@
Due to the current lack of appropriate evaluation metrics in the field of video generation and editing, automatic assessments based on CLIP often do not align well with the perceptual quality of videos. Therefore, we primarily validate the effectiveness of our method through user studies. We invited 40 volunteers to participate in our user study. The volunteers were asked to evaluate the edited videos based on three criteria: (1) Alignment with the target text prompt. (2) Inter-frame consistency of the edited video. (3) Preservation of the original video’s motion patterns and unedited regions. For each criterion, the volunteers chose the best-performing edited result from those produced by our method and the five baselines. We then calculated the proportion of times each method’s edited results were selected as the best for each criterion. As shown in Tab. 1, our method achieves comparable or superior performance across all three evaluation metrics compared to other SOTA methods.
4.4 Ablation Study
@@
We conduct ablation studies below to validate the effect of each proposed component. As shown in the first row of experimental results in Fig. 5, directly applying DDS for video editing leads to the presence of contaminations near the edited objects in the resulting video. The results in the second row show that by using ASSD instead of DDS, and smoothing the gradient information with adjacent frames, the contaminations in the resulting video are reduced, leading to improved video quality. When we further incorporate the proposed weighted attention fusion mechanism, using features from the reference branch, the contaminations are further reduced. Additionally, we observe that the edited video retains more information from the original video, such as colors and other details. Due to the significantly weaker robustness and generalization to different text conditions of current open-source text-to-video models compared to Stable Diffusion, the latter can sometimes provide higher-quality update gradients for certain text prompts, despite lacking the ability to model temporal information. As shown in the last row of Fig. 5, the updated gradients from Stable Diffusion further reduce some of the contaminations in the video and enhanced the detail quality of the edited subjects.
5 Conclusion and Limitations
@@
We propose a new score distillation-based paradigm for video editing and introduce a novel score distillation method called Adaptive Sliding Score Distillation (ASSD). ASSD mitigates errors in the gradient information provided by text-to-video models through an adaptive sliding window-based smoothing operation and adaptively extracted masks. Our pipeline also incorporates auxiliary guidance from Stable Diffusion to provide more precise gradients for updating the latent code of original video. Additionally, to better capture the information from the original video to correct errors in the edited video and further improve the quality of the update gradients, we employ proposed weighted attention fusion. This modifies certain spatial self-attention layers in pre-trained text-to-video models to support the fusion of information between the editing branch and the reference branch. Experimental results validate the effectiveness and potential of our new paradigm in video editing tasks, as well as the efficacy of our proposed components.
@@
It should be noted that the performance of our method heavily relies on the capabilities of the text-to-video models providing the gradient information for editing. The effectiveness of such methods relies on the premise that the pre-trained text-to-video model can provide an update gradient that, despite containing errors, is still effective. Given the current lack of sufficiently powerful open-source text-to-video models, the upper bound of our method’s performance is significantly constrained.

### v2

@@ Introduction
Weighted Attention Fusion
Although the proposed ASSD effectively reduces errors in the update gradients, local editing errors and over-editing still occur, which can be observed in Fig. 4. It is desirable to preserve some information in the original video, including the layout of the original video, the details in the unedited parts, and some attributes in the edited parts such as the motion pattern of the edited object. Intuitively, features from the original video can be utilized to reduce over-editing, thereby further enhancing the editing quality. To better preserve the information and leverage features in the original video, we introduce a Weighted Attention Fusion module.
Specifically, in the spatial attention layers of ModelScopeT2V or Zeroscope, the self-attention for the latent code $\bm{z}^{i}$ corresponding to the $i$-th frame in the video is:
where $\bm{W}_{q}(\cdot)$, $\bm{W}_{k}(\cdot)$, $\bm{W}_{v}(\cdot)$ are linear projections, and $d$ is the scaling factor. We use $\bm{z}^{i}$ to denote the latent code corresponding to the $i$-th frame in the editing branch of ASSD and use $\bm{\hat{z}}^{i}$ to denote the latent code corresponding to the $i$-th frame in the reference branch of ASSD. The proposed reference-attention between them is defined as:
Then, our weighted attention fusion can be implemented as:
$\mathrm{AttnFusion}_{i}=(1-\lambda)\cdot\mathrm{SelfAttn}_{i}+\lambda\cdot\mathrm{RefAttn}_{i},$ (13)
where $\lambda\in[0,1]$, it controls the balance between the extent of editing and the preservation of information from the original video. In the majority of cases, we set it to 0.8. This module allows for the fusion of features from the editing branch with those from the reference branch. We replaced the self-attention calculations in the last few spatial attention layers of the U-Net only in the editing branch with our proposed weighted attention fusion, and find this design significantly improves the quality of the edited videos.
Experiments
@@
TAV (Wu et al. 2023) 0.2988 0.9086 19.06 2.64 4.25
FateZero (Qi et al. 2023) 0.2738 0.9324 20.63 26.17 30.62
ControlVideo (Zhang et al. 2023b) 0.2916 0.9301 8.48 5.30 3.70
FLATTEN (Cong et al. 2023) 0.2931 0.9249 14.82 17.99 11.11
TokenFlow (Geyer et al. 2023) 0.2910 0.9561 12.67 20.11 24.40
@@
Experiments Setting
We collected over 30 text-video pairs from public datasets like DAVIS (Perazzi et al. 2016), WebVid (Bain et al. 2021), and YouTube (Abu-El-Haija et al. 2016), extracting 8 or 16 evenly spaced frames from each video. We used ModelscopeT2V and Zeroscope as our base models. For optimizing the latent code of one video, we employed the SGD optimizer with a learning rate of 0.1 for 200 steps, taking approximately 3 to 4 minutes on a single V100 GPU.
To evaluate the performance of our proposed method, we evaluated our method against five SOTA video editing methods: Tune-A-Video (TAV) (Wu et al. 2023) fine-tunes the denoising U-Net on a single video to learn the structural integrity of the original video. FateZero (Qi et al. 2023) achieves zero-shot video editing via DDIM Inversion and attention manipulation. FLATTEN (Cong et al. 2023) uses an external optical flow detection model for accurate attention manipulation. TokenFlow (Geyer et al. 2023) ensures consistency by modifying keyframes and propagating changes throughout the video. ControlVideo (CV) (Zhang et al. 2023b) leverages depth and edge information from the original video to control editing results.
Qualitative Experiment
Fig. 3 presents the qualitative comparison results between our method and other SOTA baselines. Our method achieves comparable or superior performance in terms of alignment with the target text prompts, preservation of the original video’s layout, motion patterns, and unedited regions, while allowing for geometric transformation of the edited objects. When an editing method allows for significant geometric transformations of the edited objects, it typically introduces challenges in preserving non-edited regions. Conversely, methods that struggle with geometric transformations of edited objects generally achieve better preservation of non-edited regions. Existing methods often face one of two issues. ControlVideo and Tune-A-Video allow for geometric transformations of the original video to better match the target prompt but fail to preserve the non-edited parts of the original video. FateZero, FLATTEN, and TokenFlow achieve better preservation of non-edited regions but struggle to support significant geometric transformations. Our method strikes a better balance between supporting geometric transformations and preserving non-edited regions.
Quantitative Experiment
@@
We utilized the CLIP model (Radford et al. 2021) for automatic evaluation, employing two metrics: CLIP Text Alignment Score and CLIP Frame Consistency Score. CLIP Text Alignment Score is defined as the mean cosine similarity between the CLIP image embeddings of the frames in the edited video and the CLIP text embedding of the target text prompt. This metric measures how well the frames in the edited video aligns with the target text prompt. CLIP Frame Consistency Score is defined as the mean cosine similarity between the CLIP image embeddings of frames within the video. This metric assesses the inter-frame consistency of the edited video. The results of the automatic metric are reported in Tab. 1. Although our method achieved only average results on both evaluation metrics, it is worth noting that there is still a lack of appropriate automatic evaluation metrics. The CLIP Frame Consistency Score focuses solely on the semantic consistency between different frames and fails to measure the continuity of motion and the consistency of objects between consecutive frames. Similarly, the CLIP Text Alignment Score only assesses the semantic alignment between the frames in the video and the target text prompt, and does not measure the preservation of non-edited parts.
Figure 5: Visualization of One-Step Prediction of $\bm{Z}_{0}$ by Different Models at Large and Small Time Steps. $\bm{Z}_{0}$ is computed from the one-step predicted noise, and since visualizing $\bm{Z}_{0}$ is more meaningful, we choose to visualize $\bm{Z}_{0}$.
Figure 6: The process of editing videos with our method. The images on the left depict the gradual transformation of a jeep into a bus, while the images on the right illustrate the gradual transformation of a jeep into a Lamborghini.
Figure 7: Ablation study for the $\lambda$ hyperparameter in Weighted Attention Fusion. By adjusting lambda, we can balance between the stability of the editing and the extent of the editing.
@@
Since automatic assessments based on CLIP often do not align well with the perceptual quality of videos, we primarily validate the effectiveness of our method through user studies. We invited 40 volunteers to participate in our user study. The volunteers were asked to evaluate the edited videos based on three criteria: (1) Alignment between the edited video and the target text prompt. (2) Inter-frame consistency and the continuity of motion between adjacent frames of the edited video. (3) Preservation of the original video’s motion patterns and unedited regions. For each criterion, the volunteers chose the best-performing edited result from those produced by our method and the five baselines. We then calculated the proportion of times each method’s edited results were selected as the best for each criterion. As shown in Tab. 1, our method achieves comparable or superior performance across all three evaluation metrics compared to other SOTA methods. Since FateZero typically only edits the color and texture of objects and struggles with geometric transformations, as seen in Fig. 3 of the qualitative experiments, it effectively facilitates the preservation of non-edited parts and object IDs, resulting in the best performance in Preservation.
Ablation Study
We conduct ablation studies to validate the effect of each proposed component and the choice of hyperparameter.
@@
As shown in the first row of experimental results in Fig. 4, directly applying DDS for video editing leads to the presence of contaminations near the edited objects in the resulting video. The results in the second row show that by using ASSD to adaptively smoothing the gradient information with adjacent frames, the contaminations in the resulting video are reduced, leading to improved video quality. When we further incorporate the proposed Weighted Attention Fusion mechanism, using features from the reference branch, the contaminations are further reduced. Additionally, we observe that the edited video retains more information from the original video, such as colors and other details. Due to the significantly weaker robustness and generalization to different text conditions of current open-source T2V diffusion models compared to Stable Diffusion, we introduced Image-based Joint Guidance as the supervision. As shown in the last row of Fig. 4, the updated gradients from Stable Diffusion further reduce some of the contaminations in the video and enhanced the detail quality of the edited subjects. In Fig. 7, we validate different choices for the $\lambda$ hyperparameter in Weighted Attention Fusion. When $\lambda$ is large, more information from the reference branch is introduced into the editing branch, reducing contaminations in the output but also weakening the degree of editing. Conversely, when $\lambda$ is small, the opposite effect is observed.
Conclusion and Limitations
@@
We introduce a novel score distillation-based paradigm for video editing, featuring Adaptive Sliding Score Distillation (ASSD). ASSD mitigates gradient errors from T2V diffusion models through adaptive sliding window-based smoothing and extracted masks. Our pipeline incorporates joint guidance from Stable Diffusion for precise latent code updates. We employ weighted attention fusion, modifying spatial self-attention layers in pre-trained T2V diffusion models to fuse information between editing and reference branches. Experimental results validate the efficacy of our paradigm and its components in video editing tasks.
@@
The efficacy of these methods hinges on the assumption that pre-trained T2V diffusion models provide effective update gradients, despite inherent errors. However, the dearth of robust open-source T2V diffusion models significantly constrains our method’s performance ceiling.